# TNO, The Hague, The Netherlands

### v1

@@ 3.3 Logistic Network Design
We now shift from network design to location assignment problems (LAP). In the first paper on this theme, Satori et al. satori2023quantum propose a novel approach for solving the LAP in an Automated Storage and Retrieval System (AS/RS) using quantum annealing. The problem involves optimising the assignment of product locations on shelves to improve picking efficiency. The proposed method considers product pairs based on picking frequency and assigns them to empty shelves in order of distance from an outlet, with the decision variable being the swapping of positions within product pairs. The paper introduces two formulations: the All-Layout formulation and the Swap formulation. The former is formulated as a mixed-integer programming (MIP) problem, while the latter is designed for iterative optimisation, reducing the number of variables. The authors evaluate the performance of both formulations using various solvers, including a MIP-solver (COIN-OR),SA, QA, and hybrid QA. Numerical experiments validate the expected retrieve time as an objective function and demonstrate the effectiveness of iterative optimisation in the Swap formulation. The paper concludes that the proposed Swap formulation, particularly when combined with quantum annealing, offers a promising approach for solving this specific problem.
@@
Next, the chapter by Guo guo2020quantum discusses the application of QA to solve, what they call, NP-hard spatial optimisation problems. Spatial optimisation involves mathematical models to find the best solutions for spatial decision problems, such as the FLP and NDP. The case study involves programming a $p-$median model using the D-Wave QA and applying quantum annealing to a biomass-to-biofuel supply chain optimisation problem. Numerical results suggest a computational advantage of quantum annealing over classical simulated annealing in certain scenarios. The authors suggest that quantum computing, particularly quantum annealing, holds promise for solving complex geospatial problems, but further research is needed to understand its full potential and integrate it with modern High Performance Computing HPC architecture. The authors envision a near-term future where QPUs could supplement HPC for specific spatial optimisation problems, with a long-term potential to be a game changer in solving computationally demanding geospatial problems.
@@
The paper by Giraldo et al. giraldo2022using contributes to the integration of quantum computing into urban and regional science, focusing on the maximal covering location problem (MCLP). The MCLP involves optimising population coverage within a specified service distance by locating a fixed number of facilities. The paper outlines the process of transforming the MCLP into a format suitable for two quantum computing paradigms: QA (D-Wave) and GBC (IBM). The authors conduct computational experiments on real quantum devices and classical quantum simulators, specifically applying QA and QAOA. The results reveal the successful solution of MCLP instances using quantum annealing technology. The paper compares the performance of QA and QAOA, noting that quantum annealing shows great potential for solving this kind of problems, while gate-based quantum technology faces (relatively more) challenges.
@@
Facility Location Problems have connections with the Quadratic Assignment Problem. The quadratic nature of this problem makes it, expectedly, suited for quantum computing. The paper by Khumalo khumalo2022investigation discusses the application of quantum computing to solve two specific problems: the TSP and the Quadratic Assignment Problem (QAP). The TSP was already discussed before. The authors explore again the potential of VQE and QAOA, in solving instances of TSP and QAP. The paper extends previous benchmarks by including the newest and largest IBM quantum devices. It introduces preliminary findings for the QAOA on tractable instances and presents metrics to assess the feasibility spectrum of quantum algorithms on IBM devices. The authors investigate the impact of the conditional reset feature on IBM systems, comparing it to previous results. The experimental results demonstrate that classical optimisation algorithms outperform quantum algorithms in terms of success rate, feasibility, and computational time for solving TSP and QAP instances for now. The VQE algorithm performs better than QAOA in terms of solution quality metrics. The conditional reset feature on IBM devices does not show a noticeable improvement in computational performance. The paper suggests that as quantum technology evolves, the performance of quantum algorithms could improve. The authors highlight the need for further exploration of quantum formulations, such as QUBO and Alternating Direction Method of Multipliers (ADMM), and the potential advancements in higher-performing processors with more qubits.
@@
Also Tosun et al. tosun2022new dive into the QAP. Their goal is to minimise the total assignment cost in a scenario with $n$ facilities and $n$ locations. This cost is calculated by summing the facility placement costs and multiplying inter-location distances with flow amounts among facilities. The paper discusses Quantum Bridge Analytics, Quantum Computing, and the mathematical formulation of QAP as a QUBO model. It explores the challenges and potential of combining metaheuristics with quantum computing for solving optimisation problems efficiently. The authors propose a methodology to transform large QAP instances into QUBO models and compare results with classical robust tabu search methods. The study suggests that even with current quantum computing limitations, QUBO transformation can be feasible for difficult instances, paving the way for hybrid techniques that combine classical and quantum computing. The authors acknowledge challenges such as finding optimal penalty values and plan to automate this process in future research.
@@
As a last example, the work of Choo et al. chooinvestigating explores the feasibility of solving the FLP as a Quadratic Assignment Problem (QAP) on D-Wave’s QA. The authors propose a method to convert the QAP into a QUBO formulation and embed it onto D-Wave’s hardware. The study reveals that, in the worst case, the current hardware can only handle QAPs with seven variables due to the limitation of qubits. To address this, the authors demonstrate a reduction in qubits required for QAPs with duplicate rows, potentially allowing the solution of larger instances. Additionally, they introduce a decomposition approach for solving generic QAPs on D-Wave’s QA. The work emphasises the challenges and limitations of directly solving QAPs on D-Wave’s hardware and provides insights into potential improvements as quantum computing technology evolves.
@@ 3.4 Fleet maintenance and optimisation
The first work in this direction is the paper by Vikstaal et al. vikstaal2020applying . They explore the application of the QAOA to instances of this problem derived from real-world data. The instances are reduced to fit on quantum devices with 8, 15, and 25 qubits, leaving only one feasible solution per instance. The reduction allows mapping the Tail Assignment problem onto the Exact Cover problem. The QAOA is simulated on an ideal quantum computer to investigate its performance in solving the simplified Tail Assignment problem. The study reveals that repeated runs of QAOA can identify feasible solutions with close to unit probability for all instances. The authors also observe patterns in the variational parameters, enabling the use of an interpolation strategy that simplifies the classical optimisation part of the QAOA. The study also highlights non-trivial properties in the connectivity of the instances and raises questions about the performance of QAOA compared to classical algorithms for larger instances of the problem. The authors also mention an alternative method for optimising variational parameters using the Gibbs objective function, leaving it for further study.
@@
Next, the dissertation of Martins martins2020applying aims to study the feasibility of solving the TAP, considering operational restrictions and costs, using Quantum Annealing. The main contributions of the dissertation include detailing a model for the Tail Assignment Problem, analysing and comparing the scalability of different modelling techniques, and presenting a comparison of solutions obtained through classical and hybrid approaches. The scalability analysis indicates that Direct QUBO modelling performs best, and the comparison between solvers suggests that Hybrid Solvers outperform others in terms of solution quality.
@@
This paper by Willsch et al. willsch2022benchmarking conducts a benchmark of the QPUs of two prominent quantum annealers: the D-Wave Advantage (with 5000+ qubits) and its precursor D-Wave 2000Q (with 2000+ qubits). One of the benchmarking problems used is the TAP. The benchmark set encompasses problems of diverse sizes, connectivity levels, and complexities. Results of the study are that the Advantage system exhibits superior performance across almost all problems, showcasing a noteworthy increase in success rates and the ability to handle larger problems compared to D-Wave 2000Q. The improved performance of Advantage is attributed not only to the increased qubit count but also to enhanced qubit connectivity. Advantage’s Pegasus topology, featuring 15 connections per qubit, surpasses D-Wave 2000Q’s Chimera topology with only 6 connections. The Advantage system demonstrates approximately twice the speed of D-Wave 2000Q in terms of programming and readout times. It also displays shorter time-to-solution and annealing times, along with consistently higher success rates and smaller fluctuations over multiple repetitions. D-Wave 2000Q may achieve better success rates for problems with sparse connectivity that do not necessitate the many new couplers present on Advantage. Improved connectivity does not universally enhance performance, depending on the specific requirements of each problem instance.
@@ 3.5 Cargo-loading, knapsack and bin-packing problems
The paper by Nayak et al. nayak2022quantum focuses on addressing challenges in the aviation industry related to the optimisation of aircraft cargo loading and the reduction of loading/unloading operations at multiple stopovers. The primary goals are to maximise the payload capacity of aircraft, leading to reduced fuel consumption and increased revenue, and to minimise the operational and handling costs associated with loading and unloading cargo. The paper presents a quantum approach for tackling these challenges, experimenting with both QA-based and GBC algorithms. Among the quantum platforms tested, D-Wave’s QA is identified as the most promising, as it offers sufficient qubits to address realistic problems compared to gate-based and quantum-inspired optimisers. The experiments are conducted on a Boeing 747-8F aircraft, one of the largest cargo carriers in the world. The model used for the experiments manages the 2D/3D stacking of containers with and without racks. Results obtained from quantum solvers, particularly D-Wave, are said to outperform classical solvers, demonstrating the potential of quantum computing in solving complex cargo loading problems. The authors emphasise that due to the limitations of qubits in current quantum hardware, their experiments had to be constrained to minimal parameters.
@@
Also, Sotelo et al. hernandez2020application ; sotelo2021determination created an application of quantum optimisation techniques to cargo logistics on ships and airplanes. Their papers address the optimisation challenge of determining the optimal configuration for maximising the loading of cargo vehicles while adhering to specific constraints: maximum volume available and maximum possible weight. The proposed solution, presented at the Airbus Quantum Computing Challenge, employs the VQE algorithm. The Airbus Quantum Computing Challenge consisted of 5 problems: 1. Aircraft Climb Optimisation, 2. Computational Fluid Dynamics, 3. Quantum Neural Networks for Solving Partial Differential Equations, 4. Wingbox Design Optimisation and last 5. Aircraft Loading Optimisation (ALO). The solution was implemented in the IBM Qiskit platform, utilising a noise-free environment simulator. The problem set comprises 200 different container arrangements for loading, with random starting points and 1000 iterations for the VQE algorithm. Realistic weight distributions are ensured by selecting weights with an expected value of 1/3 of the maximum weight of the problem instance. Results are evaluated using a metric based on the logarithm of the median of normalised energies for 200 instances. The experiments compare the performance of seven classical optimisers available in Qiskit. The findings indicate variations in the convergence speed of classical algorithms, with the Nelder-Mead algorithm demonstrating superior results. The convergence analysis reveals that over 70% of problem instances achieve a convergence ratio of 0.01 for the Nelder-Mead algorithm, while other classical algorithms exhibit varying degrees of convergence efficiency. For instance, the COBYLA algorithm achieves a 50% convergence ratio.
@@
Also the work of Traversa traversa2019aircraft and Pilon et al. pilon2021aircraft published a solution to the ALO problem as formulated by Airbus. Unlike developing algorithms for future quantum computers, the approach by Traversa formulates the ALO problem using Integer Programming (IP) and explores its solution through the MemComputing paradigm (which is not quantum). The ALO problem involves optimising the placement of containers with different sizes and weights in an aircraft, considering constraints on maximum weight, shear, and centre of gravity. The study proposes a two-objective IP problem with binary variables, utilising the MemCPU software for efficient solution strategies. The MemComputing approach, based on self-organising algebraic gates, proves to be a non-algorithmic solution with promising scaling properties, demonstrating its potential for solving large-scale ALO problems efficiently. Pilon formulates this problem using a QUBO formulation. The model’s performance is benchmarked on different solvers, including a classical solver for QUBO functions and the D-Wave 2000Q Quantum Annealer. The results indicate the feasibility of the QUBO formulation and highlight comparisons between classical and quantum approaches in terms of computational time and loaded weight.
@@
There are several papers looking directly at the Bin Packing Problem (BPP). The work of Romero et al. v2023hybrid introduces a hybrid quantum-classical framework, called Q4RealBPP, designed to solve real-world instances of the three-dimensional Bin Packing Problem (3dBPP), involving efficient packing of items into bins. The work in romero2023solving adds to that framework in the following way: i) the existence of heterogeneous bins, ii) the extension of the framework to solve not only three-dimensional, but also one- and two-dimensional instances of the problem, iii) requirements for item-bin associations, and iv) delivery priorities. Q4RealBPP addresses realistic characteristics such as package and bin dimensions, overweight restrictions, affinities among item categories, and preferences for item ordering, making it applicable to actual industrial and logistics scenarios. They successfully apply Q4RealBPP to 12 instances of different nature, showcasing its capacity to handle real-world constraints, based on (D-Wave’s) Leap Constrained Quadratic Model (CQM). They outline the need for a thorough comparison between Q4RealBPP and traditional artificial intelligence methods, considering factors such as robustness of results and execution times.
@@
Another work on the BPP is by De Andoin et al. de2022comparative ; de2022hybrid . These papers explore the potential of quantum and hybrid quantum-classical algorithms to provide advantageous solutions for the one-dimensional BPP (1dBPP). The proposed hybrid approach integrates a quantum annealing subroutine for sampling feasible solutions, followed by a classical optimisation subroutine for problem solution construction. To assess its performance, the paper compares this hybrid strategy with classical alternatives, specifically random sampling and a random-walk-based heuristic. The focus is on evaluating the quantum approach’s ability to avoid stagnation, a common issue in classical algorithms. The benchmark comprises 18 instances with 10 and 12 packages, incorporating various weight distributions. Results indicate that the quantum annealing strategy avoids stagnation observed in classical algorithms, enabling a faster full sampling of feasible partial solutions. The hybrid approach outperforms the random walk in a majority of cases. To measure the performance of the sampling strategy, a metric counting the number of iterations needed to measure the full feasible partial solution space is proposed. The paper emphasises the importance of hyperparameter optimisation, acknowledging that a more refined selection could further enhance algorithm performance.
@@
The study bu Cellini et al cellini2023qal introduces QAL-BP, a novel QUBO approach specifically for BPP. QAL-BP stands for Quantum Augmented Lagrangian approach for Bin Packing problem. It utilises an augmented Lagrangian method to incorporate the bin packing constraints into the objective function while also facilitating an analytical estimation of heuristic, but empirically robust, penalty multipliers. This approach leads to a more versatile and generalisable model that eliminates the need for empirically calculating instance-dependent Lagrangian coefficients, a requirement commonly encountered in alternative QUBO formulations for similar problems. To assess the effectiveness of the proposed approach, they conduct experiments on a set of bin-packing instances using D-Wave’s Advantage system. Additionally, they compare the results with those obtained from two different classical solvers, namely simulated annealing and Gurobi. The experimental findings confirm the correctness of the proposed formulation and also demonstrate the potential of quantum computation in effectively solving the bin-packing problem, particularly as more reliable quantum technology becomes available.
@@
The next related problem is the Knapsack Problem (KP). Van Dam et al. van2021quantum introduces enhancements to the QAOA, to make it more suitable for Knapsack Problem (KP). The KP involves selecting items with weights and values to maximise the total value under a weight constraint. The contributions of the paper are two techniques applied to QAOA. They first use the outcome of a classical greedy algorithm to define an initial quantum state and mixing operation for QAOA. Next they use quantum exploration to avoid local minima around the greedy solutions obtained from the classical algorithm. The study compares the performance of constant-depth quantum optimisation heuristics with similarly shallow classical greedy and simulated annealing algorithms. Despite the likelihood that more sophisticated classical algorithms may outperform quantum algorithms, the comparison is considered fair as it involves algorithms of equal simplicity. The paper suggests that the quantum algorithms perform well with weak dependency on instance-specific fine-tuning.
@@
The same problem, however in a slightly different application area, is studied in the work by Benson et al. benson2023cqm . This paper demonstrates the use of D-Wave’s Leap Hybrid solver for solving a KP, specifically selecting meal combinations from a fixed menu that fit certain constraints. The optimisation problem is formulated as a Constrained Quadratic Model and uses D-Wave’s CQM solver. The paper also discusses the generalisation of this model for finding optimal drug molecules within a large search space with complex and contradictory structures and property constraints. The CQM-solver is shown to solve multi-object optimisation problems on an expedited timescale, making it a valid choice for molecular drug design. While the presented problem has a manageable number of combinations, the authors discuss the scalability of quantum annealing for larger optimisation problems, highlighting potential advantages over classical computing.
@@
A variation of the KP that is even more likely to benefit from quantum computing is the Quadratic Knapsack Problem (QKP). Bontekoe et al. bontekoe2023translating recognise that, in general, constrained quadratic binary optimisation problems can be translated into QUBOs in several ways, which can have a large impact on the performance when solving the QUBO. They show six different QUBO formulations for the QKP and compare their performance using simulated annealing. The best performance is obtained by a formulation that uses no auxiliary variables for modelling the inequality constraints.

### v2

@@ 3.3 Logistic Network Design
We now shift from network design to location assignment problems (LAP). In the first paper on this theme, Satori et al. [134] propose a novel approach for solving the LAP in an Automated Storage and Retrieval System (AS/RS) using QA. The problem involves optimising the assignment of product locations on shelves to improve picking efficiency. The proposed method considers product pairs based on picking frequency and assigns them to empty shelves in order of distance from an outlet, with the decision variable being the swapping of positions within product pairs. The paper introduces two formulations: the All-Layout formulation and the Swap formulation. The former is formulated as a mixed-integer programming (MIP) problem, while the latter is designed for iterative optimisation, reducing the number of variables. The authors evaluate the performance of both formulations using various solvers, including a MIP-solver (COIN-OR),SA, QA, and hybrid QA. Numerical experiments validate the expected retrieve time as an objective function and demonstrate the effectiveness of iterative optimisation in the Swap formulation. The paper concludes that the proposed Swap formulation, particularly when combined with QA, offers a promising approach for solving this specific problem.
@@
Next, the chapter by Guo [59] discusses the application of QA to solve, what they call, NP-hard spatial optimisation problems. Spatial optimisation involves mathematical models to find the best solutions for spatial decision problems, such as the FLP and NDP. The case study involves programming a $p-$median model using the D-Wave QA and applying QA to a biomass-to-biofuel supply chain optimisation problem. Numerical results suggest a computational advantage of QA over classical simulated annealing in certain scenarios. The authors suggest that quantum computing, particularly QA, holds promise for solving complex geospatial problems, but further research is needed to understand its full potential and integrate it with modern High Performance Computing HPC architecture. The authors envision a near-term future where QPUs could supplement HPC for specific spatial optimisation problems, with a long-term potential to be a game changer in solving computationally demanding geospatial problems.
@@
The paper by Giraldo et al. [54] contributes to the integration of quantum computing into urban and regional science, focusing on the maximal covering location problem (MCLP). The MCLP involves optimising population coverage within a specified service distance by locating a fixed number of facilities. The paper outlines the process of transforming the MCLP into a format suitable for two quantum computing paradigms: QA (D-Wave) and GBC (IBM). The authors conduct computational experiments on real quantum devices and classical quantum simulators, specifically applying QA and QAOA. The results reveal the successful solution of MCLP instances using QA technology. The paper compares the performance of QA and QAOA, noting that QA shows great potential for solving this kind of problems, while gate-based quantum technology faces (relatively more) challenges.
@@
Facility Location Problems have connections with the Quadratic Assignment Problem. The quadratic nature of this problem makes it, expectedly, suited for quantum computing. The paper by Khumalo [75] discusses the application of quantum computing to solve two specific problems: the TSP and the Quadratic Assignment Problem (QAP). The TSP was already discussed before. The authors explore again the potential of VQE and QAOA, in solving instances of TSP and QAP. The paper extends previous benchmarks by including the newest and largest IBM quantum devices. It introduces preliminary findings for the QAOA on tractable instances and presents metrics to assess the feasibility spectrum of quantum algorithms on IBM devices. The authors investigate the impact of the conditional reset feature on IBM systems, comparing it to previous results. The experimental results demonstrate that classical optimisation algorithms outperform quantum algorithms in terms of success rate, feasibility, and computational time for solving TSP and QAP instances for now. The VQE algorithm performs better than QAOA in terms of solution quality metrics. The conditional reset feature on IBM devices does not show a noticeable improvement in computational performance. The paper suggests that as quantum technology evolves, the performance of quantum algorithms could improve. The authors highlight the need for further exploration of quantum formulations, such as QUBO and Alternating Direction Method of Multipliers (ADMM), and the potential advancements in higher-performing processors with more qubits.
@@
Also Tosun et al. [153] dive into the QAP. Their goal is to minimise the total assignment cost in a scenario with $n$ facilities and $n$ locations. This cost is calculated by summing the facility placement costs and multiplying inter-location distances with flow amounts among facilities. The paper discusses Quantum Bridge Analytics, Quantum Computing, and the mathematical formulation of QAP as a QUBO model. It explores the challenges and potential of combining metaheuristics with quantum computing for solving optimisation problems efficiently. The authors propose a methodology to transform large QAP instances into QUBO models and compare results with classical robust tabu search methods. The study suggests that even with current quantum computing limitations, QUBO transformation can be feasible for difficult instances, paving the way for hybrid techniques that combine classical and quantum computing. The authors acknowledge challenges such as finding optimal penalty values and plan to automate this process in future research.
@@
As a last example, the work of Choo et al. [34] explores the feasibility of solving the FLP as a Quadratic Assignment Problem (QAP) on D-Wave’s QA. The authors propose a method to convert the QAP into a QUBO formulation and embed it onto D-Wave’s hardware. The study reveals that, in the worst case, the current hardware can only handle QAPs with seven variables due to the limitation of qubits. To address this, the authors demonstrate a reduction in qubits required for QAPs with duplicate rows, potentially allowing the solution of larger instances. Additionally, they introduce a decomposition approach for solving generic QAPs on D-Wave’s QA. The work emphasises the challenges and limitations of directly solving QAPs on D-Wave’s hardware and provides insights into potential improvements as quantum computing technology evolves.
@@ 3.4 Fleet maintenance and optimisation
The first work in this direction is the paper by Vikstaal et al. [156]. They explore the application of the QAOA to instances of this problem derived from real-world data. The instances are reduced to fit on quantum devices with 8, 15, and 25 qubits, leaving only one feasible solution per instance. The reduction allows mapping the Tail Assignment problem onto the Exact Cover problem. The QAOA is simulated on an ideal quantum computer to investigate its performance in solving the simplified Tail Assignment problem. The study reveals that repeated runs of QAOA can identify feasible solutions with close to unit probability for all instances. The authors also observe patterns in the variational parameters, enabling the use of an interpolation strategy that simplifies the classical optimisation part of the QAOA. The study also highlights non-trivial properties in the connectivity of the instances and raises questions about the performance of QAOA compared to classical algorithms for larger instances of the problem. The authors also mention an alternative method for optimising variational parameters using the Gibbs objective function, leaving it for further study.
@@
Next, the dissertation of Martins [94] aims to study the feasibility of solving the TAP, considering operational restrictions and costs, using QA. The main contributions of the dissertation include detailing a model for the Tail Assignment Problem, analysing and comparing the scalability of different modelling techniques, and presenting a comparison of solutions obtained through classical and hybrid approaches. The scalability analysis indicates that Direct QUBO modelling performs best, and the comparison between solvers suggests that Hybrid Solvers outperform others in terms of solution quality.
@@
This paper by Willsch et al. [161] conducts a benchmark of the QPUs of two prominent quantum annealers: the D-Wave Advantage (with 5000+ qubits) and its precursor D-Wave 2000Q (with 2000+ qubits). One of the benchmarking problems used is the TAP. The benchmark set encompasses problems of diverse sizes, connectivity levels, and complexities. Results of the study are that the Advantage system exhibits superior performance across almost all problems, showcasing a noteworthy increase in success rates and the ability to handle larger problems compared to D-Wave 2000Q. The improved performance of Advantage is attributed not only to the increased qubit count but also to enhanced qubit connectivity. Advantage’s Pegasus topology, featuring 15 connections per qubit, surpasses D-Wave 2000Q’s Chimera topology with only 6 connections. The Advantage system demonstrates approximately twice the speed of D-Wave 2000Q in terms of programming and readout times. It also displays shorter time-to-solution and annealing times, along with consistently higher success rates and smaller fluctuations over multiple repetitions. D-Wave 2000Q may achieve better success rates for problems with sparse connectivity that do not necessitate the many new couplers present on Advantage. Improved connectivity does not universally enhance performance, depending on the specific requirements of each problem instance.
@@ 3.5 Cargo-loading, knapsack and bin-packing problems
The paper by Nayak et al. [103] focuses on addressing challenges in the aviation industry related to the optimisation of aircraft cargo loading and the reduction of loading/unloading operations at multiple stopovers. The primary goals are to maximise the payload capacity of aircraft, leading to reduced fuel consumption and increased revenue, and to minimise the operational and handling costs associated with loading and unloading cargo. The paper presents a quantum approach for tackling these challenges, experimenting with both QA-based and GBC algorithms. Among the quantum platforms tested, D-Wave’s QA is identified as the most promising, as it offers sufficient qubits to address realistic problems compared to gate-based and quantum-inspired optimisers. The experiments are conducted on a Boeing 747-8F aircraft, one of the largest cargo carriers in the world. The model used for the experiments manages the 2D/3D stacking of containers with and without racks. Results obtained from quantum solvers, particularly D-Wave, are said to outperform classical solvers, demonstrating the potential of quantum computing in solving complex cargo loading problems. The authors emphasise that due to the limitations of qubits in current quantum hardware, their experiments had to be constrained to minimal parameters.
@@
Also, Sotelo et al. [64, 145] created an application of quantum optimisation techniques to cargo logistics on ships and airplanes. Their papers address the optimisation challenge of determining the optimal configuration for maximising the loading of cargo vehicles while adhering to specific constraints: maximum volume available and maximum possible weight. The proposed solution, presented at the Airbus Quantum Computing Challenge, employs the VQE algorithm. The Airbus Quantum Computing Challenge consisted of 5 problems: 1. Aircraft Climb Optimisation, 2. Computational Fluid Dynamics, 3. Quantum Neural Networks for Solving Partial Differential Equations, 4. Wingbox Design Optimisation and last 5. Aircraft Loading Optimisation (ALO). The solution was implemented in the IBM Qiskit platform, utilising a noise-free environment simulator. The problem set comprises 200 different container arrangements for loading, with random starting points and 1000 iterations for the VQE algorithm. Realistic weight distributions are ensured by selecting weights with an expected value of 1/3 of the maximum weight of the problem instance. Results are evaluated using a metric based on the logarithm of the median of normalised energies for 200 instances. The experiments compare the performance of seven classical optimisers available in Qiskit. The findings indicate variations in the convergence speed of classical algorithms, with the Nelder-Mead algorithm demonstrating superior results. The convergence analysis reveals that over 70% of problem instances achieve a convergence ratio of 0.01 for the Nelder-Mead algorithm, while other classical algorithms exhibit varying degrees of convergence efficiency. For instance, the COBYLA algorithm achieves a 50% convergence ratio.
@@
Also the work of Traversa [154] and Pilon et al. [116] published a solution to the ALO problem as formulated by Airbus. Unlike developing algorithms for future quantum computers, the approach by Traversa formulates the ALO problem using Integer Programming (IP) and explores its solution through the MemComputing paradigm (which is not quantum). The ALO problem involves optimising the placement of containers with different sizes and weights in an aircraft, considering constraints on maximum weight, shear, and centre of gravity. The study proposes a two-objective IP problem with binary variables, utilising the MemCPU software for efficient solution strategies. The MemComputing approach, based on self-organising algebraic gates, proves to be a non-algorithmic solution with promising scaling properties, demonstrating its potential for solving large-scale ALO problems efficiently. Pilon formulates this problem using a QUBO formulation. The model’s performance is benchmarked on different solvers, including a classical solver for QUBO functions and the D-Wave 2000Q Quantum Annealer. The results indicate the feasibility of the QUBO formulation and highlight comparisons between classical and quantum approaches in terms of computational time and loaded weight.
@@
There are several papers looking directly at the Bin Packing Problem (BPP). The work of Romero et al. [124] introduces a hybrid quantum-classical framework, called Q4RealBPP, designed to solve real-world instances of the three-dimensional Bin Packing Problem (3dBPP), involving efficient packing of items into bins. The work in [123] adds to that framework in the following way: i) the existence of heterogeneous bins, ii) the extension of the framework to solve not only three-dimensional, but also one- and two-dimensional instances of the problem, iii) requirements for item-bin associations, and iv) delivery priorities. Q4RealBPP addresses realistic characteristics such as package and bin dimensions, overweight restrictions, affinities among item categories, and preferences for item ordering, making it applicable to actual industrial and logistics scenarios. They successfully apply Q4RealBPP to 12 instances of different nature, showcasing its capacity to handle real-world constraints, based on (D-Wave’s) Leap Constrained Quadratic Model (CQM). They outline the need for a thorough comparison between Q4RealBPP and traditional artificial intelligence methods, considering factors such as robustness of results and execution times.
@@
Another work on the BPP is by De Andoin et al. [8, 9]. These papers explore the potential of quantum and hybrid quantum-classical algorithms to provide advantageous solutions for the one-dimensional BPP (1dBPP). The proposed hybrid approach integrates a QA subroutine for sampling feasible solutions, followed by a classical optimisation subroutine for problem solution construction. To assess its performance, the paper compares this hybrid strategy with classical alternatives, specifically random sampling and a random-walk-based heuristic. The focus is on evaluating the quantum approach’s ability to avoid stagnation, a common issue in classical algorithms. The benchmark comprises 18 instances with 10 and 12 packages, incorporating various weight distributions. Results indicate that the QA strategy avoids stagnation observed in classical algorithms, enabling a faster full sampling of feasible partial solutions. The hybrid approach outperforms the random walk in a majority of cases. To measure the performance of the sampling strategy, a metric counting the number of iterations needed to measure the full feasible partial solution space is proposed. The paper emphasises the importance of hyperparameter optimisation, acknowledging that a more refined selection could further enhance algorithm performance.
@@
The study bu Cellini et al [30, 31] introduces QAL-BP, a novel QUBO approach specifically for BPP. QAL-BP stands for Quantum Augmented Lagrangian approach for Bin Packing problem. It utilises an augmented Lagrangian method to incorporate the bin packing constraints into the objective function while also facilitating an analytical estimation of heuristic, but empirically robust, penalty multipliers. This approach leads to a more versatile and generalisable model that eliminates the need for empirically calculating instance-dependent Lagrangian coefficients, a requirement commonly encountered in alternative QUBO formulations for similar problems. To assess the effectiveness of the proposed approach, they conduct experiments on a set of bin-packing instances using D-Wave’s Advantage system. Additionally, they compare the results with those obtained from two different classical solvers, namely simulated annealing and Gurobi. The experimental findings confirm the correctness of the proposed formulation and also demonstrate the potential of quantum computation in effectively solving the bin-packing problem, particularly as more reliable quantum technology becomes available.
Also Gatti et al. [53] use a BPP formulation to tackle aircraft load optimization. Using QA, they maximizes load characteristics like weight and volume while ensuring flight stability. Results from D-Wave simulations show promising improvements in efficiency and scalability for real-world cargo aircraft models.
Last, Matt et al. [96] introduce a QAOA approach for solving a BPP-alike problem, the irregular strip packing problem, optimizing material usage in industries. The approach decomposes the problem into two sub-problems: the TSP and rectangular packing. Using quantum optimization algorithms like QAOA and QOBA, the algorithm minimizes waste and improves efficiency, outperforming classical methods in experiments.
@@
The next related problem is the Knapsack Problem (KP). Van Dam et al. [40] introduces enhancements to the QAOA, to make it more suitable for Knapsack Problem (KP). The KP involves selecting items with weights and values to maximise the total value under a weight constraint. The contributions of the paper are two techniques applied to QAOA. They first use the outcome of a classical greedy algorithm to define an initial quantum state and mixing operation for QAOA. Next they use quantum exploration to avoid local minima around the greedy solutions obtained from the classical algorithm. The study compares the performance of constant-depth quantum optimisation heuristics with similarly shallow classical greedy and simulated annealing algorithms. Despite the likelihood that more sophisticated classical algorithms may outperform quantum algorithms, the comparison is considered fair as it involves algorithms of equal simplicity. The paper suggests that the quantum algorithms perform well with weak dependency on instance-specific fine-tuning. Christiansen et al. [35] integrate the Quantum Tree Generator (QTG) with a Grover-mixer QAOA framework and present a new method called Amplitude Amplification-mixer QAOA (AAM-QAOA) for solving the KP. Experimental results show that AAM-QAOA outperforms the state-of-the-art Copula-QAOA for smaller instances, but both methods struggle for larger instances at lower circuit depths. However, for sufficiently high circuit depths, AAM-QAOA can eventually sample the optimal solution, offering promising results for optimizing combinatorial problems like the knapsack problem.
@@
The same problem, however in a slightly different application area, is studied in the work by Benson et al. [15]. This paper demonstrates the use of D-Wave’s Leap Hybrid solver for solving a KP, specifically selecting meal combinations from a fixed menu that fit certain constraints. The optimisation problem is formulated as a Constrained Quadratic Model and uses D-Wave’s CQM solver. The paper also discusses the generalisation of this model for finding optimal drug molecules within a large search space with complex and contradictory structures and property constraints. The CQM-solver is shown to solve multi-object optimisation problems on an expedited timescale, making it a valid choice for molecular drug design. While the presented problem has a manageable number of combinations, the authors discuss the scalability of QA for larger optimisation problems, highlighting potential advantages over classical computing. Also Bozejko et al. [21] use D-Wave’s quantum machine and compute lower and upper bounds, enhancing efficiency and solution quality for KP problems in complex production systems.
@@
A variation of the KP that is even more likely to benefit from quantum computing is the Quadratic Knapsack Problem (QKP). Bontekoe et al. [18] recognise that, in general, constrained quadratic binary optimisation problems can be translated into QUBOs in several ways, which can have a large impact on the performance when solving the QUBO. They show six different QUBO formulations for the QKP and compare their performance using simulated annealing. The best performance is obtained by a formulation that uses no auxiliary variables for modelling the inequality constraints.
Different approaches can be found the next works. In Shirai et al. [141] a postprocessing variationally scheduled quantum algorithm (pVSQA) for constrained combinatorial optimization problems is proposed. It combines variational methods and postprocessing techniques to improve solution quality and efficiency. Applied to graph partitioning and quadratic knapsack problems, pVSQA demonstrates superior performance on both simulators and real quantum devices. Ardelean et al. [10] introduce a hybrid quantum search algorithm with genetic optimization (HQAGO) to solve complex KPs efficiently. By fixing some qubits as classical bits, it reduces the search space and computational complexity. Lastly, Cui at al. [38] propose a hybrid quantum search algorithm for the Multi-Dimensional Knapsack Problem (MDKP). It combines Grover’s search with classical preprocessing and iterative refinement, reducing depth, width, and runtime. The method shows improved efficiency and scalability compared to existing quantum and classical algorithms.